OpenStudy (anonymous):

82percent of what number is 73.8

OpenStudy (kewlgeek555):

Here's whatcha do: \[(73.8\div82)\times100\]

OpenStudy (nincompoop):

82/100 = 73.8/x solve for x get the reciprocal of both sides 100/82 = x/73.8 then multiply both sides by 73.8

OpenStudy (cakt07):

73.8 /82 X 100

OpenStudy (cakt07):

NP

OpenStudy (kewlgeek555):

You divide 73.8 into 82 parts because you know it is 82%, then you get the answer and multiply it by 100 to get the full 100%.

OpenStudy (nincompoop):

while you arrive at the same answer with the methods the other people suggested, the more systematic way of doing it is the solution I have indicated.